# Photonic entanglement enhanced multidimensional spectroscopy for probing exciton correlations: theory and applications to photosynthetic aggregates

### v1

@@ Appendix A Notes on the Hamiltonian
The exciton nonlinearity values (absolute magnitudes), $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ used in the construction of the two-exciton Hamiltonian, i.e., $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}+U_{m}^{(1)}\delta_{mk}\delta_{nl}\delta_{mn}+U_{mn}^{(2)}\delta_{mk}\delta_{nl}$ are displayed below in Fig. 5:
Figure 5: The absolute exciton nonlinearity values are listed here in the order in which they appear in the local two-exciton basis. For a detailed discussion, see section II.1 and section III.2.3.
Only the values of $U_{m}^{(1)}$ differ between the two plots.
@@ Appendix B Outline of derivation of the signal
Here, we succinctly outline the derivation, directing the reader to [16, 49, 22] for a more involved discussion. The DQC signal is defined as the integrated rate of change of the transmitted intensity of the detected photon mode (e.g., denoted $E_{4}$), expressed as $S=\int_{-\infty}^{\infty}dt_{4}\langle E_{4}^{\dagger}(t_{4})E_{4}\rangle$. Here, the averaging is performed over the time-dependent exciton-photon density operator. Expanding the density operator to the third power of the exciton-photon interaction, and introducing the the frequency domain variables, we obtain:
@@ Appendix C Notes on the Joint spectral intensity function
In Fig. 6, we present the squared Joint Spectral Amplitude (JSA) for better clarity. An increase in the entanglement time parameter is inversely correlated with the number of significant singular values. As a consequence, the parametric freedom to choose the anti-correlated frequency pairs is diminished.
Figure 6: The squared Joint Spectral Amplitude (JSA) is presented for two distinct entanglement time parameters: $\tilde{T}_{\text{ent}}=10$ fs (left column) and $60$ fs (right column). The lower panels display the distribution of the 2020 largest normalized singular values.

### v2

@@ Appendix A Notes on the Hamiltonian
Figure 5: The absolute values of the exciton nonlinearity parameters are plotted. The sequence follows the order in which the values appear in the local two-exciton basis. Values in the upper panel are used in the simulations for Fig. 2 and Fig. 3, while those in the lower panel are used in Fig. 4. Further details regarding the Hamiltonian and these results are provided in Section II.1 and  III.2.3.
The exciton Hamiltonian contains two types of exciton nonlinearity parameters, namely $U_{m}^{(1)}$ and $U_{mn}^{(2)}$. They affect overtone and combination two-exciton terms in the Hamiltonian, via $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}+U_{m}^{(1)}+U_{mn}^{(2)}$. In this work, we use two sets of exciton nonlinearity parameters that differ by the magnitudes of the overtone nonlinearities only. The absolute magnitudes of these two sets are plotted in Fig. 5.
@@ Appendix B Outline of derivation of the signal
In this appendix, we succinctly outline the derivation, directing the reader to [16, 49, 22] for a more involved discussion. The DQC signal is defined as the integrated rate of change of the transmitted intensity of the detected photon mode, denoted $E_{4}$. It is expressed as $S=\int_{-\infty}^{\infty}dt_{4}\langle E_{4}^{\dagger}(t_{4})E_{4}\rangle$. Here, the averaging is performed over the time-dependent joint exciton-photon density operator which involves partial trace over the phonon modes. Expanding the density operator to the third power of the exciton-photon interaction, and introducing the the frequency domain variables, we obtain:
@@ Appendix C Notes on the Joint spectral intensity function
For a better understanding of the correlation properties of the entangled photon pairs, in Fig. 6, we plot the squared Joint Spectral Amplitude (JSA) and the singular values obtained via singular value decomposition of the kernel. An increase in the entanglement time parameter is inversely correlated with the number of statistically significant singular values. Consequently, the parametric freedom to select anti-correlated frequency pairs is diminished.
Figure 6: The squared Joint Spectral Amplitude is presented for two distinct entanglement time parameters: $\tilde{T}_{\text{ent}}=10$ fs (left column) and $60$ fs (right column). Two lower panels display the distribution of the 2020 largest normalized singular values.